x+y+z=8
x-y-z=-4
-x+y-z=-8
Please solve this system of equations.

Answer by chitra(359)

The given system of equations are:

x + y + z = 8 -----------(1)

x - y - z = -4 --------- (2)

-x + y - z = -8 -----------(3)


Adding (1) and (2), we get:

2x = 4


this implies x = 2


Adding (2) and (3), we get:

-2z = -12

z = 6

Substituting the values of x and zin (1), we get y = 0

Thus the solution.

x = 2, y = 0 and z = 6